HUBBLE SPACE TELESCOPE DAILY REPORT #916 PERIOD COVERED: 0800 EST Daily Status Report as of 176/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4277 (Calibration of Nearby Galaxy Type Ia Supernovae as Standard Candles) The Wide Field Camera was used to make two images of Cepheid variable stars located inside galaxy NGC-5253. This will make it possible to accurately determine the distance to this galaxy. This distance will be used to find the true luminosity of the type - Ia supernova which have been observed in this galaxy. Both observations were received and displayed in the OSS. The observations Were reported to be okay. 1.2 Completed WFC 5010 (Determination of Extragalactic Distance Scale. III. M-101. Recovery) The Wide Field Camera was used to make two images of the galaxy M-101 located in Ursa Major looking for Cepheid variable stars. This will make it possible to accurately determine the distance to this galaxy. Both observations were reported to look fine. 1.3 Completed HSP/POL 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make four calibration observations of standard star VID998 to determine the instruments long term sensitivity of the IDT detectors. All four observations were received and displayed. The first observations showed the image of the target slightly off center. The next observation left it well centered. The following calibration images were reported to be good. 1.4 Completed Two Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make snapshot images of the protoplanetary system PK118+08D1 to study its current morphology. Two Sets of observations were received and displayed. There was no visible target in either set of images. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed FOS/BL 3483 (Lyman Alpha Emission in Galaxies: The Case of Galaxies Causing the Z = 0) The Faint Object Spectrograph was used to make Lyman Alpha emission line observations of the galaxy PKS-1127-1 which is near the line of sight of a known quasi-stellar object. This galaxy is expected to produce Magnesium II absorption lines in the QSOs spectra. This can be used to confirm that star formation is taking place. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 7 Successful : 7 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Note: Due to an operational oversight the NSSC-1 status buffer was dumped today while the FOS was transmitting science data. This led to an Science Data Formatter (SDF)lockout condition which was corrected by real-time command procedures approximately one hour later at 176/1315Z. Science data was lost for two FOS Observations. The HST Operations prepared to make the transition to the new PRS (PORTS Refurbishment System) 2.7 system this afternoon (Friday, June 25, 1993) at 2:00 P.M. The new system will be running with a hot backup for the remainder of this weekend. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 3483 (Lyman Alpha Emission in Galaxies: the Case of Galaxies Causing the Z = 0) 3.2 Complete Two Sets of WFC 4277 (Calibration fo Nearby Galaxy Type Ia Supernovae as Standard Candles) 3.3 Complete FGS 4031 (High Speed Astrometry - A Search for Planetary Companions to Low Mass) 3.4 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.5 Complete Two Sets of GHRS/2 4067 (GHRS 1ST Order Grating spectral Lamp Test) 3.6 Complete Two Set of FOC/96 3801 (Search for QSOs Suitable for Subsequent Observation of He II) 3.7 Complete PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) 3.8 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.9 Complete Three Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies With FOC/96) 3.10 Complete GHRS/2 4926 (GHRS SSA Pickup Test - Part 2) 3.11 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.12 Complete FOS/RD 4213 (Spectrophotometry of the Lensed "Clover Leaf" Broad Absorption line) 3.13 Complete HSP/UV2 4209 (HSP Detector Sensitivity Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-